What is the price of a bond (to the nearest cent) with 6 years to maturity, 5.9% coupon rate, semiannual payments, par of $1000, and the yield to maturity of 4.97%? SHOW WORK WITH ALL STEPS
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

Please find the detailed answer as follows:

Semi-Annual Interest Payment = 1000*5.9%*1/2 = 28.5

Semi-Annual Rate of Interest = 4.97%/2 = 2.485%

Current Bond Price = 29.5/(1+.02485)^1 + 29.5/(1+.02485)^2 + 29.5/(1+.02485)^3 + 29.5/(1+.02485)^4 + 29.5/(1+.02485)^5 + 29.5/(1+.02485)^6 + 29.5/(1+.02485)^7 + 29.5/(1+.02485)^8 + 29.5/(1+.02485)^9 + 29.5/(1+.02485)^10 + 29.5/(1+.02485)^11 + 29.5/(1+.02485)^12 + 1000/(1+.02485)^12 = $1047.74

Answer is $1047.74